# The Unreasonable Ineffectiveness of the Deeper Layers

### v1

@@ Abstract
We empirically study a simple layer-pruning strategy for popular families of open-weight pretrained LLMs, finding minimal degradation of performance on different question-answering benchmarks until after a large fraction (up to half) of the layers are removed. To prune these models, we identify the optimal block of layers to prune by considering similarity across layers; then, to “heal” the damage, we perform a small amount of finetuning. In particular, we use parameter-efficient finetuning (PEFT) methods, specifically quantization and Low Rank Adapters (QLoRA), such that each of our experiments can be performed on a single A100 GPU. From a practical perspective, these results suggest that layer pruning methods can complement other PEFT strategies to further reduce computational resources of finetuning on the one hand, and can improve the memory and latency of inference on the other hand. From a scientific perspective, the robustness of these LLMs to the deletion of layers implies either that current pretraining methods are not properly leveraging the parameters in the deeper layers of the network or that the shallow layers play a critical role in storing knowledge.
@@ 1 Introduction
Over the last few years, large language models (LLMs) have evolved from mere research artifacts Radford et al. (2019) into useful products OpenAI (2022). To a large extent this evolution can be attributed to a dramatic increase in scale of the resources used for training OpenAI (2023); Gemini Team et al. (2023). Since these models will likely see most of their total lifetime FLOPs in inference mode after training completes, the pretraining of LLMs requires not only considerations for efficient, i.e. compute-optimal, training Kaplan et al. (2020); Hoffmann et al. (2022), but also requires inference awareness De Vries (2023); Sardana and Frankle (2023).
What about models that have already been trained? Beyond the training considerations indicated by neural scaling laws, there are also numerous post-training techniques for reducing the cost and time of finetuning and then inferencing LLMs. In particular, quantization can be used to reduce the memory footprint of models by decreasing the precision of the model weights Dettmers et al. (2022); Frantar et al. (2022); Dettmers and Zettlemoyer (2023); Xiao et al. (2023), Low Rank Adapters (LoRA) can be used to reduce the cost of finetuning and customization by only updating a small subset of the model parameters Hu et al. (2021), or pruning can be used to reduce the memory footprint and time for inference by directly eliminating unnecessary parameters or connections LeCun et al. (1989); Hassibi and Stork (1992); Han et al. (2015); Li et al. (2016); Frankle and Carbin (2018). As these three strategies are more or less orthogonal, in a resource constrained environment ideally we would want to leverage all three of these post-training efficiency techniques in combination. Towards that direction, the popular method of QLoRA Dettmers et al. (2023) introduced a handful of innovations that enabled 4-bit quantization of parameters and LoRA finetuning to work together.
@@
Building on that combination, in this work we study a very simple pruning strategy using open-weight LLMs. In particular, we develop a method that uses the similarity between the representations at different layers to identify the optimal layers to prune for a given pruning fraction; then, after removing these layers we “heal” the pruning-induced mismatch with a small amount of fine tuning (using QLoRA). Our main result is that we can remove a substantial fraction of the deepest layers from models with minimal degradation in downstream performance. For example, for Llama-2-70B Touvron et al. (2023a) we can eliminate up to roughly half of the layers before the performance collapses. An overview of our strategy and the results of pruning Llama-2-70B are shown in Figure 1.
@@
Figure 1: Overview of our layer-pruning strategy and example results: (a) a flowchart describing the algorithm: if removing $n$ layers, we find the layer, $\ell^{*}$, that minimizes the angular distance, $d$, between layers $\ell$ and $\ell\!+\!n$; we then remove the $n$ layers beginning with layer $\ell^{*}$; finally, if necessary, we can “heal” the damage with a small amount of (parameter-efficient) finetuning. (b) a schematic depicting the removal of $n$ total layers, indexed from $\ell^{*}\!$ to $\ell^{*}\!\!+\!n\!-\!1$. (c) angular distance, $d$, between different numbers of layers, $n$, vs. the layer number, $\ell$, that indexes the beginning of the block of $n$; the bottom curve (darkest purple) represents $n=1$, while the top curve (lightest yellow) represents $n=64$; the black line traces $\ell^{*}(n)$, the minimum of the angular distance across the different sized layer blocks. (d) results of pruning Llama-2-70B with healing (dark blue) and without healing (light blue) as a function of the fraction of layers removed: the top (middle) panel gives the accuracy on the MMLU (BoolQ) question-answering benchmark, while the bottom panel the autoregressive loss on a subset of the C4 validation set; here, the dashed red lines (dashed gray lines) indicate the accuracy or loss of the original unpruned model (of random guessing); these plots illustrate that typical behavior we find in which there are sharp transitions in performance for the accuracy of question-answering tasks (here between 40%-50% pruning fraction), but continuity and very slow growth in the healed loss (dark blue) up to at least to 80% pruning fraction.
@@
Pruning is not only useful for reducing the footprint of inference, but also for understanding how a network uses its parameters: if you can remove large blocks of the network with minimal effect on its performance, then those blocks are likely not very important. In particular, our intuition for dropping layers comes considering the residual structure of the transformer architecture. In more detail, the output of the final layer can be decomposed as a sum over the outputs of all the model layers plus the embedded input. If such a sum had numerous and independent terms, then removing a handful of them should not significantly change the output. However, since the terms are not independent – each layer is input to the following layer – we should expect to be able to remove terms if the residual contribution from a particular layer is small. In other words, if the output of each layer does not change too much from layer to layer. [^1] This is strongly suggested by “lens” investigations that studied the evolution of the token distribution as a function of layer index such as the “logit lens” nostalgebraist (2020) and the “tuned lens” Belrose et al. (2023). A separate line of reasoning along these lines previously inspired neural ODEs Chen et al. (2018), and led Ref. Yang et al. (2023) to argue that ideally representation should change substantially from layer to layer in order to most effectively make use of the parameters of a network.
@@
In conjunction with our layer pruning, we investigate the similarity of layer representations at different separations and find broadly that deeper layers are qualitatively more similar to neighboring layers than shallow layers (with the exception of the very final layer). This suggests an even simpler pruning strategy: remove layers beginning at the penultimate layer and proceed from deep to shallow until the desired number of layers have been removed. [^2] This strategy is especially interesting in situations where resource constraints inhibit the full application of the similarity-informed pruning algorithm described in Figure 2(a). In this case, we find that, after healing the damage with a small amount of QLoRA finetuning, we can achieve performance that nearly matches the more involved similarity-informed layer pruning strategy. The effectiveness of this method is evidence that LLMs might not properly leverage the parameters in the deeper layers of the network.
Overall, we hope you take these three bulleted points with you:
  • The model’s memory footprint and inference time decreases linearly with the number of removed layers. [^3] Contrast this with quantization: the memory footprint decreases with the quantization ratio, but the inference time remains approximately fixed since parameters are typically de-quantized before any FLOPs. This makes layer pruning a powerful tool, especially if the model’s performance is robust to dropping layers.
  • All the efficiency methods – pruning, PEFT and quantization – can be effectively combined with each other. Thus, in this work each experiment was performed on a single A100 GPU and is easily accessible to the open source and academic communities.
  • The robustness of models to removing the deeper layers, the sharp transition in performance on downstream knowledge tasks (e.g. MMLU and BoolQ), and the smooth behavior of the autoregressive loss with respect to those pruning fractions, altogether suggest that the shallow layers may play a critical role in storing knowledge.
@@
The structure of this paper is as follows. In §2, we first perform a literature review of both practical post-training strategies and science-of-deep-learning investigations that motivate our work. Then, in §3, we give intuition for our layer pruning strategy and explain our method in detail, while in §4 we iterate over all our experimental results. Finally, we conclude in §5 by highlighting directions of future work. Specific model, finetuning, dataset, and evaluation details can be found in Appendix A, and evaluations ablations can be found in Appendix B.
Note: As we were finalizing this work, a preprint of Ref. Men et al. (2024) was posted, which has a number of points of overlap with our work.
@@ 2 Literature Review
In this section, we review practical strategies for post-training efficiency and discuss some scientific investigations that provide motivation for, or insight into, our approach: in §2.1, we first review the history of pruning and then discuss its modern application to LLMs; in §2.2, we contrast pruning with distillation, an alternative strategy for reducing the parameter count of LLMs; then in §2.3, we discuss the various practical methods for efficient finetuning and inference acceleration that can be used in conjunction with our pruning strategy; finally in §2.4 we highlight some scientific investigations into some depth-dependent statistical properties of LLMs that are complementary to our results.
2.1 Pruning
Pruning is a method for reducing the size of a trained machine-learning model by removing unnecessary parameters, either individually or together as a group. Pruning for neural networks has a long history (LeCun et al., 1989; Hassibi and Stork, 1992), and, as originally conceived, unstructured pruning techniques sparsify networks by removing individual parameters based on pre-defined criteria. For instance, if a parameter of the model has a very small value, then removing it – i.e. by setting it to exactly zero – will likely have minimal impact on performance. Inspired by this early work, modern researchers began exploring different criteria for such unstructured pruning, focusing mostly on computer vision models (Han et al., 2015; Chen et al., 2015; Srinivas and Babu, 2015). In particular, Ref. Han et al. (2015) developed an iterative pruning method for alternatively pruning and finetuning a network in order to reach better compression ratios and performance.
While these models were smaller, they were not necessarily more efficient: sparsifying networks by removing individual parameters according to a criterion leads to irregular or pseudorandom sparsification patterns that are difficult to accelerate without specialized hardware or libraries designed for sparsity Li et al. (2016). To that end, structured pruning techniques were developed to remove irrelevant groups of parameters together, such as particular channels or filters in convolutional networks. As this increased their practical relevance, researchers then began exploring structured pruning across computer vision (Li et al., 2016; Wen et al., 2016; Hu et al., 2016; He et al., 2017; Huang et al., 2018) and pre-transformer NLP architectures (Murray and Chiang, 2015; See et al., 2016; Kim and Rush, 2016).
Following unprecedented progress in language modeling, recent work has focused on applying structured pruning methods to the Transformer Vaswani et al. (2017). These studies consider nearly every possible component of the model architecture for elimination, with methods ranging from dropping attention heads (Voita et al., 2019; Michel et al., 2019; Kim and Awadalla, 2020), to dropping layers (Fan et al., 2019; Zhang and He, 2020; Fan et al., 2021; Jha et al., 2023; Sajjad et al., 2023; Liu et al., 2023a), to pruning hidden states (Hou et al., 2020), to rank reducing large weight matrices Sharma et al. (2023), replacing sparse weight matrices with smaller dense ones Ashkboos et al. (2024), to many combinations of the aforementioned groups (Xia et al., 2022; Lagunas et al., 2021).
Of the prior work that also considers transformer layer dropping, most Fan et al. (2019); Zhang and He (2020); Fan et al. (2021); Xia et al. (2022); Sajjad et al. (2023) study BERT-style models Devlin et al. (2018), while we consider decoder-only GPT-style models Radford et al. (2019) that are most commonly used for large-scale language modeling and generation. BERT-style models are naturally suited for understanding tasks due to their bidirectional masked language modeling (MLM) objective, while GPT-style models are instead suited for generation, due to their autoregressive objective. While this divide has been questioned in light of more powerful GPT-style models (Zhong et al., 2023), previous work (Ethayarajh, 2019) has found significant qualitative differences between BERT and GPT models in terms of the evolution of the layer-wise representation of words. Altogether, this suggests that layer-dropping strategies will behave differently between the two families.
One study for BERT-style pre-trained models, Ref. Sajjad et al. (2023), concludes that the best layer-pruning strategy is dropping the final layers; this partially resonates with our results, although in contrast we find that (a) for some pruning sizes keeping the last few layers of the model is actually beneficial, and that (b) for all pruning sizes keeping the very last layer is essential. Additionally, while the authors also study similarity between representations in different layers – as in our approach – they actually found a higher similarity between representations in the shallow layers compared to the deeper ones – which very sharply disagrees with our results. Importantly, the models considered in Ref. Sajjad et al. (2023) consist of a few hundred million parameters, which is much smaller than the model scales we consider in our work. Perhaps as a consequence, the authors didn’t observe the sharp transition in downstream accuracies that we report in §4.1, despite the fact that they also finetuned their pruned models.
In contrast, while Ref. Jha et al. (2023) does consider GPT-style models, the methodology is quite different from ours: (i) rather than pretraining first and then using a fixed layer-dropping strategy as we do, instead the authors incrementally drop layers in a modified pretraining procedure; and (ii) the authors study their own sub-1B parameter models, while we focus on the families of readily available, open-weight, large-scale 2.7B-70B parameter models that are commonly used and/or finetuned for practical applications.
Finally, a systematic approach to layer dropping in transformers has also been studied in the context of wav2vec models, which are encoder-only models that map speech to embeddings and are sized in the hundred-million parameter regime Baevski et al. (2020). With these models, Ref. Liu et al. (2023a) developed a layer-pruning algorithm based on the correlation between layers and downstream metrics. Beyond the model architecture and domain, one significant difference between this and our work is that Ref. Liu et al. (2023a) considered non-contiguous pruning proposals, e.g. dropping alternate layers. Our intuition for layer pruning predicts that this shouldn’t work as well – at least for decoder-only language models – as it creates multiple mismatches, one with each block of layers removed.
2.2 Model distillation
A completely different method for reducing the size of a trained machine-learning model is model distillation Hinton et al. (2015), in which knowledge is transferred from a large “teacher” model to a smaller “student” model by training the student on the distribution predicted by the teacher. The essential insight is that this can transform the very general knowledge and capabilities of the teacher into more streamlined, compressed, and possibly skill-specific representations.
While a very general technique, in the setting of language models, distillation has been implemented with (a) white-box approaches, in which the the student is trained to imitate the teacher’s logits Gu et al. (2023) or hidden states Jiao et al. (2019); as well as with (b) black-box approaches, in which the student only has access to the output tokens generated by the teacher. This latter approach broadly covers cases where the student is trained on text that is augmented by the teacher in some way, such as by adding synthetic labels Wang et al. (2021), generating high quality synthetic text Eldan and Li (2023); Li et al. (2023a); Gunasekar et al. (2023) by providing chain of thought reasoning Fu et al. (2023); Hsieh et al. (2023), which aims to enhance the student’s reasoning skills, or by annotating instructions that enhance the student’s instruction-following capabilities Jiang et al. (2023a).
2.3 Efficient finetuning and inference acceleration
Complementary to directly reducing size of a model, parameter-efficient finetuning (PEFT) focuses on reducing the cost of specializing LLMs to certain tasks. In particular, Low Rank Adapters (LoRA) reduce the memory and compute of fine tuning by freezing the pretrained model and introducing a parametrically small number of additional trainable weights Hu et al. (2021). We use its quantized cousin, QLoRA Dettmers et al. (2023), to keep our experiments cost efficient. Other PEFT methods that can be combined with our work are Refs. Li et al. (2023b) and Zhang et al. (2023): in the first, the initialization of the LoRA matrices is adjusted to a quantization scheme; in the second, LoRA ranks for different LLM modules are chosen in an adaptive manner.
For additional efficiency gains we could combine our layer-pruned models with methods that further accelerate inference: with speculative decoding Leviathan et al. (2023), tokens are rapidly generated from a smaller draft model and then evaluated in parallel by the main model; with Medusa Cai et al. (2024) the draft model is discarded for extra decoding heads, but ultimately achieves a similar effect. In particular, it could be interesting to consider highly-compressed layer-pruned models as potential draft models in a speculative decoding setup.
2.4 A breadth of depth-dependent studies
Finally, let us highlight some scientific work that study the depth-dependent properties of LLMs. One relevant direction considers how knowledge and linguistic properties are encoded in language models. On the one hand, Refs. Meng et al. (2022); Dai et al. (2021) analyze the storage and recall of factual associations: these works emphasize that knowledge localizes within the middle Meng et al. (2022) or final Dai et al. (2021) layers, which has implications for directly editing or erasing part of a model’s factual knowledge. On the other hand, attempts to perform such editing gives evidence that information may be stored non-locally across layers Hase et al. (2023). Relatedly, Ref. Geva et al. (2023) investigates the way facts are processed during inference, distinguishing between the role of attention heads, for attribute extraction, and the MLP blocks, for subject enrichment: both are delocalized across several layers.
Next, following the earlier “logic lens” nostalgebraist (2020), Ref. Belrose et al. (2023) invented a technique they called “tuned lens” to study the trajectory of predictions by using a learnable affine transformation to convert intermediate representations into a distributions over tokens (see also Din et al. (2023)). By studying the layer-to-layer dynamics of this distribution, the authors noted that it tended to converge. This convergence is very suggestive that that the deeper layers could be prunable, while the fact that they had to train an affine probe is likely related to our observation that the final layer cannot be pruned. Somewhat relatedly, Ref. Gurnee and Tegmark (2023) observed that geographic features in the underlying text can be determined from linear probes trained on intermediate activations, as long as the activations are deeper than halfway.
More abstractly, Refs. Voita et al. (2023); Liu et al. (2023b) found that the sparsity of activations transitions at around halfway through a network’s forward pass, evolving from sparse to dense. Perhaps relatedly, Ref. Panigrahi et al. (2023) investigated which model weights update the most during finetuning, finding that it’s those in the mid-layers.
@@ 3.2 Layer-pruning algorithm(s)
[…]
@@ 4 Results
For our experiments, we pruned a wide variety of large-scale LLMs from 2.7B to 70B parameters spanning 32 to 80 total unpruned layers. Specifically, we used models in the Llama-2 family Touvron et al. (2023a), the Qwen family Bai et al. (2023), Mistral-7B Jiang et al. (2023b), and Phi-2 Javaheripi and Bubeck (2023). For these models, we executed the “healing” step using QLoRA Dettmers et al. (2023): our models were quantized to 4-bit precision and then finetuned, using QLoRA for efficient training, on either 164M or 328M tokens from the Colossal Clean Crawled Corpus (C4) Raffel et al. (2020), a common pretraining dataset. As a result, each experiment of ours was performed on a single A100100100 GPU. For our QA evals, we used Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2020), a common world-knowledge and problem solving benchmark, and BoolQ Clark et al. (2019), a common yes/no reading comprehension benchmark where the answer has to be inferred from the text itself. The specifics of our models, healing procedure, dataset choices, and evaluation details can be found across Appendix A; ablations of different hyperparameter choices can be found across Appendix B.

### v2

@@ Abstract
How is knowledge stored in an LLM’s weights? We study this via layer pruning: if removing a certain layer does not affect model performance in common question-answering benchmarks, then the weights in that layer are not necessary for storing the knowledge needed to answer those questions. To find these unnecessary parameters, we identify the optimal block of layers to prune by considering similarity across layers; then, to “heal” the damage, we perform a small amount of finetuning. Surprisingly, with this method we find minimal degradation of performance until after a large fraction (up to half) of the layers are removed for some common open-weight models. From a scientific perspective, the robustness of these LLMs to the deletion of layers implies either that current pretraining methods are not properly leveraging the parameters in the deeper layers of the network or that the shallow layers play a critical role in storing knowledge. For our study, we use parameter-efficient finetuning (PEFT) methods, specifically quantization and Low Rank Adapters (QLoRA), such that each of our experiments can be performed on a single 40GB A100 GPU.
@@ 1 Introduction
In this work we study a very simple pruning strategy using open-weight LLMs. In particular, we develop a method that uses the similarity between the representations at different layers to identify the optimal layers to prune for a given pruning fraction; then, after removing these layers we “heal” the pruning-induced mismatch with a small amount of fine tuning (using QLoRA). Our main result is that we can remove a substantial fraction of the deepest layers from models with minimal degradation in downstream question-answering benchmarks. For example, for Llama-2-70B (Touvron et al., 2023) we can eliminate up to roughly half of the layers before the performance collapses. An overview of our strategy and the results of pruning Llama-2-70B are shown in Figure 1.
@@
Figure 1: Overview of our layer-pruning strategy and example results: (a) a flowchart describing the algorithm: if removing $n$ layers, we find the layer, $\ell^{*}$, that minimizes the angular distance, $d$, between layers $\ell$ and $\ell\!+\!n$; we then remove the $n$ layers beginning with layer $\ell^{*}$; finally, if necessary, we can “heal” the damage with a small amount of (parameter-efficient) finetuning. (b) a schematic depicting the removal of $n$ total layers, indexed from $\ell^{*}\!$ to $\ell^{*}\!\!+\!n\!-\!1$. (c) angular distance, $d$, between different numbers of layers, $n$, vs. the layer number, $\ell$, that indexes the beginning of the block of $n$; the bottom curve (darkest purple) represents $n=1$, while the top curve (lightest yellow) represents $n=64$; the black line traces $\ell^{*}(n)$, the minimum of the angular distance across the different sized layer blocks. (d) results of pruning Llama-2-70B with healing (light blue) and without healing (dark blue) as a function of the fraction of layers removed: the top (middle) panel gives the accuracy on the MMLU (BoolQ) question-answering benchmark, while the bottom panel the autoregressive loss on a subset of the C4 validation set; here, the dashed red lines (dashed gray lines) indicate the accuracy or loss of the original unpruned model (of random guessing); these plots illustrate that typical behavior we find in which there are sharp transitions in performance for the accuracy of question-answering tasks (here between 40%-50% pruning fraction), but continuity and very slow growth in the healed loss (light blue) up to at least to 80% pruning fraction.
@@
Our intuition for dropping layers comes from considering the residual structure of the transformer architecture. In more detail, the output of the final layer can be decomposed as a sum over the outputs of all the model layers plus the embedded input. If such a sum had numerous and independent terms, then removing a handful of them should not significantly change the output. However, since the terms are not independent – each layer is input to the following layer – we should expect to be able to remove terms if the residual contribution from a particular layer is small. In other words, if the output of each layer does not change too much from layer to layer. [^1] This is strongly suggested by “lens” investigations that studied the evolution of the token distribution as a function of layer index such as the “logit lens” (nostalgebraist, 2020) and the “tuned lens” (Belrose et al., 2023). A separate line of reasoning along these lines previously inspired neural ODEs (Chen et al., 2018), and led Yang et al. (2023) to argue that ideally representation should change substantially from layer to layer in order to most effectively make use of the parameters of a network.
@@
In conjunction with our layer pruning, we investigate the similarity of layer representations at different separations and find broadly that deeper layers are qualitatively more similar to neighboring layers than shallow layers (with the exception of the very final layer). This suggests an even simpler pruning strategy: remove layers beginning at the penultimate layer and proceed from deep to shallow until the desired number of layers have been removed. In this case, we find that, after healing the damage with a small amount of QLoRA finetuning, we can achieve performance that nearly matches the more involved similarity-informed layer pruning strategy. The effectiveness of this method is evidence that LLMs might not properly leverage the parameters in the deeper layers of the network.
That said, while question-answering (QA) benchmarks such as MMLU and BoolQ are robust to a large amount of layer pruning, other measures of performance are not: if we look at the loss on next-token predictions for an IID dataset (C4 validation set), we find that the model is smoothly damaged in proportion to the fraction of the number of layers pruned. Since perplexity typically correlates strongly with downstream metrics, this naturally begs the question: which tasks are less robust than QA benchmarks to pruning? As part of our final discussion, we explore reasoning related tasks (GSM8k and HellaSwag) and see that they are harmed by any amount of pruning. Altogether, this leads to the following accounting of state: the shallow layers likely play a critical role in the storing of knowledge and retrieving of information, while the deeper layers are important for higher-level computations such as mathematical reasoning.
@@
The structure of this paper is as follows. In §2, we first perform a literature review of both practical post-training strategies and science-of-deep-learning investigations that motivate our work. Then, in §3, we give intuition for our layer pruning strategy and explain our method in detail, while in §4 we iterate over all our experimental results. Finally, we conclude in §5 by exploring tasks beyond QA benchmarks, such as reasoning, and highlighting directions of future work. Specific model, finetuning, dataset, and evaluation details can be found in Appendix B, and evaluation ablations can be found in Appendix C.
@@ 2 Literature Review
Pruning for neural networks has a long history (LeCun et al., 1989, Hassibi and Stork, 1992): while initial work focused on unstructured pruning (Han et al., 2015, Chen et al., 2015, Srinivas and Babu, 2015), structured pruning techniques were developed to make sparse networks more efficient (Li et al., 2016, Wen et al., 2016, Hu et al., 2016, He et al., 2017, Huang et al., 2018, Murray and Chiang, 2015, See et al., 2016, Kim and Rush, 2016). Recent work, of course, focused on structured pruning of transformers (Voita et al., 2019, Michel et al., 2019, Kim and Awadalla, 2020, Fan et al., 2019, Zhang and He, 2020, Fan et al., 2021, Jha et al., 2023, Sajjad et al., 2023, Liu et al., 2023a, Hou et al., 2020, Sharma et al., 2023, Ashkboos et al., 2024, Xia et al., 2022, Lagunas et al., 2021, Men et al., 2024). Our work focuses on pruning the layers of decoder-only GPT style open-weight large language models after they’ve been pretrained. For an extended literature review, please see Appendix A.
@@ 3.2 Layer-pruning algorithm(s)
Finally, we also investigated an even simpler pruning strategy inspired by analyzing the angular distances across different model families: drop the deepest layers, excluding the final layer before the LLM head, and then (non-optionally) heal the damage. For complete clarity, this means that if we are pruning $n$ layers from an $L$-layer model, then we would remove layers $(L-n)$ to $(L-1)$, inclusive.
@@ 4 Results
For our experiments, we pruned a wide variety of large-scale LLMs from 2.7B to 70B parameters spanning 32 to 80 total unpruned layers. Specifically, we used models in the Llama-2 family (Touvron et al., 2023), the Qwen family (Bai et al., 2023), Mistral-7B (Jiang et al., 2023a), and Phi-2 (Javaheripi and Bubeck, 2023). For these models, we executed the “healing” step using QLoRA (Dettmers et al., 2023): our models were quantized to 4-bit precision and then finetuned, using QLoRA for efficient training, on either 164M or 328M tokens from the Colossal Clean Crawled Corpus (C4) (Raffel et al., 2020), a common pretraining dataset. As a result, each experiment of ours can be performed on a single 40GB A100100100100 GPU. For our QA evals, we used Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2020), a common world-knowledge and problem solving benchmark, and BoolQ (Clark et al., 2019), a common yes/no reading comprehension benchmark where the answer has to be inferred from the text itself. The specifics of our models, healing procedure, dataset choices, and evaluation details can be found across Appendix B; ablations of different hyperparameter choices can be found across Appendix C.